# VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models

### v3

@@ Appendix A Appendix
A.1 Machine Requirements for Experiments
@@
A.2 Assuring Statistical Significance of the Results within Budget for proprietary models
@@
Figure 4: Elo ratings of the model with gpt-4o judge on the full set of Arena-Hard-Auto Li et al. (2024) prompts. VARCO Arena result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category). Error bars are 95% confidence intervals.
A.3 Elo ratings from VARCO Arena compared to Human Annotations
Figure 4 shows the Elo ratings computed out of VARCO Arena. For judge, we used gpt-4o. As mentioned in the caption, the Elo ratings are bootstrapped median value from 500 trials. 95% confidence intervals also plotted as an error bar, which look negligible in scale compared to observed values. Matches are performed over Arena-Hard-Auto benchmark dataset (500 prompts).
A.4 Binary search vs. Win rate over reference
@@ A.3 Elo ratings from VARCO Arena compared to Human Annotations
3:$n_{\text{comparisons}}\leftarrow\lfloor\log_{2}(|L|)\rfloor$
4:$n_{\text{matches}}\leftarrow\lfloor|X|/n_{\text{comparisons}}\rfloor$
5:function BinarySearchPlacement($L,m_{\text{new}}$)
@@
11:         mid $\leftarrow\lfloor(\text{low}+\text{high})/2\rfloor$
12:         wins $\leftarrow 0$
@@
16:                  wins $\leftarrow$ wins $+1$
@@
31:function UpdateLeaderboard($L,m_{\text{new}}$)
32:     position, istie $\leftarrow$ BinarySearchPlacement($L,m_{\text{new}}$)
@@
A.4.1 Binary Search
We tried binary search placement of a newly added LLM to the leaderboard without reference text in Table 5. Details of how we implemented binary search are attached in Appendix 2. It turns out that binary search based on leaderboard ranks is not as reliable as the current approach of scoring the newcomer to the reference outputs. The number of judge operations performed is equivalent to the matches allocated to the least-performant model in a tournament, which is $|X|$ (i.e. maximum possible matches that an LLM could have is $|X|*\text{log}_{2}n_{\text{model}}$). Within the size of the benchmark prompts ($|X|$), binary search is incompatible with the current approach of using reference instead.
A.4.2 Comparing to the most performant Model so far: Converting Elo Table back to Win Rate
Assuming we preserved a set of match results and model outputs from the last benchmarking, we could benefit from those to perform insertion. One could pick an appropriate anchor LLM as a reference in a leaderboard to estimate the skill of a newcomer. Using previous matches from the tournaments that built the leaderboard could be used for estimating win rates over the reference. This is the same as converting the Elo table into a win rate leaderboard. Since the leaderboard is not built with full-grid matches but with tournaments, there would be some missing matches against the reference regardless we have picked. There are two ways to estimate the win rate over the reference model. We could just count the matches given are enough in amount, or we could also convert Elo ratings back to $P(i>a)$ to use it directly for scoring for the model ranks in the leaderboard. Reminding that Elo rating is purposed for expecting a likely outcome of the match, this should work. After this win rate of the newcomer model $P^{*}(n>a)=\frac{\text{count(n wins)}}{|X|}$ could be directly compared for enlisting.
@@ A.4.2 Comparing to the most performant Model so far: Converting Elo Table back to Win Rate
A.5 Separability In terms of Confidence Interval
@@
A.6 Judge configuration
A.6.1 Evaluation Prompt

### v4

@@ Appendix A Appendix
A.1 Arena-Lite Web Demo
We provide screenshots of Arena-Lite web demo here (link to Huggingface Space). You could try or locally host Arena-Lite according to its documentation. It is quite easy to use and do not require much of resources to host.
Arena-Lite provides the benchmark result (Figure 4) with helpful visualization interface that enables walking through the matches and tournaments one by one (Figure 5) and match statistics between LLMs (Figure 6). We also provide visualization that helps examining potential bias of LLM Judge being used (Figure 7).
Figure 4: Arena-Lite web screenshot 1: At the top of the result page, one can see the leaderboard of LLMs with their BT preference. If the benchmark dataset has subcategories, radar chart (right) is also visible.
Figure 5: Arena-Lite web screenshot 2: User can walk through the matches and tournaments one by one. Match brackets is visualized briefly with text UI and user can select any specific match to see the details (e.g. match result, prompt, and model outputs).
Figure 6: Arena-Lite web screenshot 3: User can see the match statistics between LLMs (i.e. win rate between model pairs, number of matches per pair and per model).
Figure 7: Arena-Lite web screenshot 4: User can see the LLM Judge’s examine how biased the LLM judge being used. The demo provides clues for potential bias toward response length and position.
A.1.1 Starter Prompt Set
We provide several judge prompts that we have used for specific target tasks. Some of those are for quite specialized tasks, and some might work for evaluating general instruction following. You could customize your own judge prompt based on your evaluation needs according to the documentation. The list of the judge prompts we provide is as follows, and one could see the detailed prompts at yaml files here.
  1. 1.
    llmbar prompt (Figure A.8.2) and conciser version of the prompt, llmbar_brief. Those are for evaluating instruction following.
  2. 2.
    translation_pair prompt for selecting translation models trained on game-specialized parallel corpora,
  3. 3.
    rag_pair_kr prompt for evaluating knowledge groundedness of korean RAG models over chatting scenario,
  4. 4.
    translation_fortunecookie prompt which was crafted for evaluating translation models specialized for translating fortune-tellings, and
  5. 5.
    post_edit prompt for evaluating conversation revision based on given persona of a speakers given in a instruction, which is quite specialized use case.
A.2 Full table for Experiment 2
Here is the extended results of Experiment 2 (Section 4.3) presented in Table 3. Aligned LLMs smaller than 7B parameters struggles to work as a proper Judge. Otherwise, Arena-Lite method excels over common practice of using encore outputs as baselines.
Table 4: Extended results for comparing Arena-Lite to baseline method of using encore outputs. We tested other LLMs as judge over various size of benchmark datasets.
Dataset size method claude 3.5 sonnet llama3.1-8b-it qwen2.5-7b-it qwen2.5-0.5b-it gemma2-2b-it
50 baseline-mediated .896 .656 .492 .010 .064
Arena-Lite (ours) .897 .715 .544 -0.051 -0.088
100 baseline-mediated .912 .732 .596 .002 .079
Arena-Lite (ours) .918 .780 .656 -0.068 -0.090
250 baseline-mediated .924 .801 .700 .045 .560
Arena-Lite (ours) .929 .830 .760 -0.131 .551
475 baseline-mediated .924 .819 .708 .083 .112
Arena-Lite (ours) .930 .845 .810 -0.131 -0.009
500 baseline-mediated .924 .820 .756 .089 .592
Arena-Lite (ours) .930 .850 .811 -0.124 .551
A.3 Machine Requirements for Experiments
@@
A.4 Assuring Statistical Significance of the Results within Budget for proprietary models
@@
Figure 8: BT preference of the model with gpt-4o judge on the full set of Arena-Hard-Auto Li et al. (2024) prompts. Arena-Lite result (bootstrapped median over 1000 samples of 500 trials) is in blue, plotted alongside the ratings from the ground truth leaderboard in red (Chatbot Arena, Hard prompts category). Error bars are 95% confidence intervals.
A.5 BT preference from Arena-Lite compared to Human Annotations
Figure 8 shows the BT preference computed out of Arena-Lite. For judge, we used gpt-4o. As mentioned in the caption, the BT preference are bootstrapped median value from 500 trials. 95% confidence intervals also plotted as an error bar, which look negligible in scale compared to observed values. Matches are performed over Arena-Hard-Auto benchmark dataset (500 prompts).
A.6 Binary search vs. Win rate over baseline
@@ A.5 BT preference from Arena-Lite compared to Human Annotations
[…]
@@
[…]
12:   wins 0$\leftarrow$ 0← 0
@@
16:      wins $\leftarrow$ wins +1+1+ 1
@@
[…]
@@
A.6.1 Binary Search
We tried binary search placement of a newly added LLM to the leaderboard without baseline output in Table 6. Details of how we implemented binary search are attached in Algorithm 2, Appendix. It turns out that binary search based on already built leaderboard ranks is not as reliable compared to utilizing the best model’s outputs as a baseline. Therefore, when adding a newcomer LLM to pre-existent leaderboard, we could utilize the already submitted responses as a baseline from the 1st placed LLM.
binary search (4o) 0.92 1.84 2.13 1.72
comp. to 1st (4o) 1.98 1.55 1.57 1.39
binary search (4o-mini) 1.27 1.82 1.21 1.5
comp. to 1st (4o-mini) 1.00 1.43 1.43 1.37
Table 5: Comparison of the binary search method versus using the top-performing model’s response as a baseline (comp. to 1st) for inserting a new LLM into the leaderboard. We report the mean rank deviation ($\bar{|\Delta_{\text{rank}}|}$) from the ground-truth leaderboard as an additional error metric. For further details, see Algorithm 2 in Appendix.
A.6.2 Comparing to the most Performant Model so far: Converting Ratings Table back to Win Rates
Assuming we preserved a set of match results and model outputs from the last benchmarking, we could benefit from those to perform insertion. One could pick an appropriate anchor LLM as a baseline in a leaderboard to estimate the skill of a newcomer. Using previous matches from the tournaments that built the leaderboard could be used for estimating win rates over the baseline. This is the same as converting the preference ratings table into a win rate leaderboard. Since the leaderboard is not built with full-grid matches but with tournaments, there would be some missing matches against the baseline regardless we have picked. There are two ways to estimate the win rate over the baseline model. We could just count the matches given are enough in amount, or we could also convert BT preference back to $P(i>a)$ to use it directly for scoring for the model ranks in the leaderboard. Reminding that BT preference rating is for expecting a likely outcome of the match, this should work. After this win rate of the newcomer model $P^{*}(n>a)=\frac{\text{count(n wins)}}{|X|}$ could be directly compared for enlisting.
@@ A.6.2 Comparing to the most Performant Model so far: Converting Ratings Table back to Win Rates
A.7 Separability In terms of Confidence Interval
@@
A.8 Judge configuration
A.8.1 Evaluation Prompt